OpenStudy (anonymous):

Solve 7 + 2(4 – x) = 7 – 2(x – 4). Answer x = 12 x = 4 This is an identity and is true for all real numbers. There is no solution.

OpenStudy (anonymous):

@dpgreen

OpenStudy (anonymous):

@iNeedTheHelp @Chlorophyll

OpenStudy (anonymous):

7+8-2x=7-2x+8 add 2x 7+8=7+8 15=15 d. all real numbers is the answer

OpenStudy (anonymous):

wait that is c.